A rock is suspended from a string; and it accelerates downward. Which one of the following statements concerning the tension in the string is true? A rock is suspended from a string; and it accelerates upward. Which statement is true concerning the tension in the string? A rock is suspended from a string and moves downward at constant speed. Which statement is true concerning the tension in the string if air resistance is ignored?

A) The tension points downward.

B) The tension is less than the weight of the rock.

C) The tension is equal to the weight of the rock.

D) The tension is greater than the weight of the rock. The tension is independent of the magnitude of the rock's acceleration.
Answer #1

Solution : when the rock is moving down wards with acceleration a then

mg- T = ma

So T = mg - ma

So tension is less than the weight .

When rope acceleration is upwards then

T - mg = ma

So T = mg +ma

It is more than the weight of rock.

When rock moves up or down without any acceleration means with constant speed than

T = WEIGHT